22. For the following equation of the path of a rocket h ( t ) = − 9.8 t 2 + 78.4 t + 196, where height, h, is measured in meters and time, t, is measured in seconds, what is the height from which it was launched (t=0)?

23.For the following equation of the path of a rocket h ( t ) = − 9.8 t 2 + 78.4 t + 196, where height, h, is measured in meters and time, t, is measured in seconds, how long after launch does it hit the ground?

24.For the following equation of the path of a rocket h ( t ) = − 9.8 t 2 + 78.4 t + 196, where height, h, is measured in meters and time, t, is measured in seconds, how long after launch does it reach its maximum height?

25.For the following equation of the path of a rocket h ( t ) = − 9.8 t 2 + 78.4 t + 196, where height, h, is measured in meters and time, t, is measured in seconds, what is the maximum height?

26.For the following equation of the path of a rocket h ( t ) = − 9.8 t 2 + 78.4 t + 196, where height, h, is measured in meters and time, t, is measured in seconds, how long after launch does it reach 196 meters?

View keyboard shortcuts12pt

Paragraph
Answer #1

Equation of the path of rocket hit -9.842 +78.4€ + 196 _22) hught when tzo is h(o)= -9.8(0)²+ 78.4(0) + 196 = 0+0+ 196 =196 =(25) Maximum height is at t=4secs h (4) = -9.8(4) ²+ 78.4(4) + 196 = -9.8 (16) + 78-4 (41 +196 - 156.8+ 313.6 + 196 = 352.8 m